OpenStudy (anonymous):

What is the arc measure of an arc with length 4.189 cm and radius equal to 3 cm. A. 480 B. 53 C. 82 D. 80

OpenStudy (anonymous):

L = r* theta where theta = arc measure theta = L / r = 4.189 / 3 radians multiply this by 180/pi gives the arc measure in degrees this comes to 80 degrres D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw